OpenStudy (anonymous):

Convert the following credit card rate to the APR. Oregon, 1 3/4% per month

OpenStudy (anonymous):

1.75%*12=21%

OpenStudy (slaaibak):

isnt it (1 + 0.0175)^12 = i + 1 then i = 23.14% ?

OpenStudy (anonymous):

Never heard of that equation. I figured it's 1.75% per month...so 12 months times 1.75% would equal the annual percentage rate.

OpenStudy (slaaibak):

well, if it's compound interest, you cant multiply it by 12, since your interests earns interest